# HAROOD: A Benchmark for Out-of-distribution Generalization in Sensor-based Human Activity Recognition

### v2

@@ 4.2. Four OOD Scenarios
We consider four real-world domain-shift scenarios: cross-person, cross-position, cross-dataset, and cross-time generalization. Each scenario reflects a common source of variability (differences across individuals, sensor locations, data sources, or temporal segments) that can change the data distribution. As shown in Table 3, these four scenarios encompass the majority of settings considered in the existing literature. To simulate these shifts, we preprocess the raw time-series (sliding-window segmentation and normalization) and split the data into separate domains accordingly. The following description details each scenario and the dataset-specific processing steps. Detailed statistics can be found in Table 4 and illustrative diagrams can be found in Figure LABEL:fig:frame-scenari. Detailed description can be found in section A.2.
@@ 5.1. Algorithms
We include not only most of the methods implemented in DomainBed (gulrajanisearch) but also several algorithms specifically designed for OOD generalization in the activity-recognition field (lu2022local; qin2023generalizable). In addition, we have organized these algorithms into categories; the detailed classification results are presented in Table 6. More details on these algorithms can be found in section A.3.
@@ 6.1. Findings
Figure 7(a) summarizes the results of HAROOD 444More results can be found in section B.2. For each task, we evaluated multiple algorithms, model-selection criteria, and model architectures. Specifically, we report the average out‑of‑distribution test accuracies and ranks over all combinations of dataset, algorithm, and model-selection criterion included in the initial release of HAROOD. These experiments encompass sixteen widely used OOD algorithms under identical conditions. The evaluational process is repeated over three independent runs using the same fixed hyperparameter set, and results are reported as mean ± standard error. From results, we draw some major conclusions:
@@ 6.2. Visualization Study
To better evaluate how well different methods extract features across tasks, we performed t-SNE visualizations on CNN-derived feature representations for both training and testing data, making comparisons across various algorithms and tasks. Figure 7 shows the visualizations of various algorithms and more results can be found in section B.5. Generally, better-performing algorithms tend to produce clearer visualizations, while poorer performance often corresponds to less distinguishable patterns.
@@ 7. Conclusion and Future Work
Appendix A More Details of HAROOD

### v3

@@ 4.1. Datasets
The six datasets used are DSADS, USC‑HAD, UCI‑HAR, PAMAP2, EMG, and WESAD. Table 2 summarizes their main statistics: number of subjects, activities, sampling rate, the number of sensor types, and approximate total instances.
@@ 4.2. Four OOD Scenarios
We consider four real-world domain-shift scenarios: cross-person, cross-position, cross-dataset, and cross-time generalization. Each scenario reflects a common source of variability (differences across individuals, sensor locations, data sources, or temporal segments) that can change the data distribution. As shown in Table 3, these four scenarios encompass the majority of settings considered in the existing literature. To simulate these shifts, we preprocess the raw time-series (sliding-window segmentation and normalization) and split the data into separate domains accordingly. The following description details each scenario and the dataset-specific processing steps. Detailed statistics can be found in Table 4 and illustrative diagrams can be found in Figure LABEL:fig:frame-scenari.
@@ 5.1. Algorithms
We include not only most of the methods implemented in DomainBed (gulrajanisearch) but also several algorithms specifically designed for OOD generalization in the activity-recognition field (lu2022local; qin2023generalizable). In addition, we have organized these algorithms into categories; the detailed classification results are presented in Table 6.
@@ 6.1. Findings
Figure 7(a) summarizes the results of HAROOD. For each task, we evaluated multiple algorithms, model-selection criteria, and model architectures. Specifically, we report the average out‑of‑distribution test accuracies and ranks over all combinations of dataset, algorithm, and model-selection criterion included in the initial release of HAROOD. These experiments encompass sixteen widely used OOD algorithms under identical conditions. The evaluational process is repeated over three independent runs using the same fixed hyperparameter set, and results are reported as mean ± standard error. From results, we draw some major conclusions:
@@
Sensitivity to window length varies across methods
Large-model baselines do not yet outperform compact HAR models
| Target | 0 | 2 |
| HARGPT | 5.48 | 5.54 |
| ERM | 67.08 | 64.79 |
We focused primarily on classic deep-learning baselines because many on-device scenarios still rely on compact models. In addition, small models can serve as encoders for larger systems. We added a large-model testing interface and initial LLM-style experiments (HARGPT (ji2024hargpt) on GPT-OSS:120b (openai2025gptoss120bgptoss20bmodel)). As shown in Table 8, the performance of large-model methods is much lower than that of small-model methods.
Model capacity and architecture jointly affect generalizatio
Inference efficiency is comparable across benchmark methods
We conducted 100 inference runs on Domain 0 of the EMG dataset to measure the inference time, and the results are reported in Table 10. Overall, the inference time across different benchmark methods remains at a comparable level, and no substantial gaps are observed among them, indicating similar computational efficiency.
F1-score largely agrees with accuracy-based comparisons
Following DomainBed, we initially adopted accuracy as the primary evaluation metric for consistency with prior studies. To provide a more comprehensive evaluation, we further report the F1-score in Table 11. The results show that the relative performance trends measured by F1-score are largely consistent with those observed using accuracy, suggesting that the comparisons among benchmark methods are stable across different evaluation metrics.
@@ 6.2. Visualization Study
To better evaluate how well different methods extract features across tasks, we performed t-SNE visualizations on feature representations for both training and testing data, making comparisons across various algorithms and tasks. Figure 7 shows the visualizations of various algorithms. Generally, better-performing algorithms tend to produce clearer visualizations, while poorer performance often corresponds to less distinguishable patterns.
@@ 7. Conclusion and Future Work
Acknowledgements.
This work was partially supported by William & Mary Faculty Research Award and Modal Academic Compute grant. The authors acknowledge William & Mary Research Computing for providing computational resources and/or technical support that have contributed to the results reported in this paper. URL: https://www.wm.edu/it/rc.Appendix A More Details of HAROOD